OpenStudy (anonymous):

Find a rational zero of the polynomial function and use it to find all the zeros of the function. f(x) = x4 + 3x3 - 5x2 - 9x - 2

OpenStudy (anonymous):

possible answers: http://prntscr.com/kvsbz

OpenStudy (amistre64):

this might be easier to do if you try some of the answer choices

OpenStudy (amistre64):

notice that there are a few differences in the options that would either make or break an option; 1, 2, and 3 if any of these makes a zero in the poly, then we know that option is correct, since none of the others contain them. If none of these provide a zero for us, then it has to be the remaining option

OpenStudy (amistre64):

do we get a zero when x=1?

OpenStudy (anonymous):

No, you get a 1... but I don't know what answer it is. >.<

OpenStudy (amistre64):

well, we have two more numbers to test out

OpenStudy (amistre64):

the one that works is the correct answer :)

OpenStudy (anonymous):

Alright.

OpenStudy (amistre64):

f(1) = 1^4 + 3(1)^3 - 5(1)^2 - 9(1) - 2 1 + 3 - 5 - 9 - 2 4 - 16 not equal 0 ; so x=1 is out

OpenStudy (amistre64):

can you work this one for me? f(2) = 2^4 + 3(2)^3 - 5(2)^2 - 9(2) - 2

OpenStudy (anonymous):

so that means it's not A?

OpenStudy (amistre64):

correct, it is not A

OpenStudy (anonymous):

f(2)=0 ?

OpenStudy (amistre64):

yes it does :) that means that our solution set HAS to have a +2 in it

OpenStudy (anonymous):

so it's d?

OpenStudy (amistre64):

i would say it is d yes; lets verify that with the wolf to be sure :)

OpenStudy (amistre64):

the wolf like it :)

OpenStudy (anonymous):

:)

OpenStudy (amistre64):

whenever i take a test that is multiple choice; then i know that one of the answers has to be correct; its just a matter of finding the most efficient way of weeding it out

OpenStudy (anonymous):

http://prntscr.com/kvvan how do I do this one?